Calculate the increase in volume when 1500cm3 of steel is heated from 0°c to 40°c take coefficient of linear expansivity as [1.2x10-5 °c-1]


1
Expert's answer
2020-10-05T10:52:24-0400

We can find the increase in volume from the formula:


"\dfrac{\Delta V}{V} = \alpha_V \Delta T,"

here, "\Delta V" is the increase in volume, "V = 1500 \ cm^3" is the volume of the steel, "\alpha_V = 3 \alpha_L" is the coefficient of volume expansivity, "\alpha_L = 1.2 \cdot 10^{-5} \ ^{\circ}C^{-1}" is the coefficient of linear expansivity and "\Delta T" is the temperature change.

Then, from this formula we can calculate the increase in volume:


"\Delta V = 3\alpha_L V \Delta T,""\Delta V = 3 \cdot 1.2 \cdot 10^{-5} \ ^{\circ}C^{-1} \cdot 1500 \ cm^3 \cdot 40 \ ^{\circ}C^{-1} = 2.16 \ cm^3,""\Delta V =2.16 \ cm^3 \cdot (\dfrac{1 \ m}{100 \ cm})^3 = 2.16 \cdot 10^{-6} \ m^3."

Answer:

"\Delta V = 2.16 \ cm^3 = 2.16 \cdot 10^{-6} \ m^3."